# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v5

@@ Abstract
We address the applicability of hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments. All the requirements posed by the operator of the network are modeled. The computational results demonstrate the readiness for application and benefits of quantum-classical hybrid solvers in the realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
@@ 1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policies, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems. Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [18]. Examples of such disturbances include late train departures/arrivals, extended dwell times, etc. Their duration can be from several minutes up to hours. A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too [5]. In the present work, we focus on train rescheduling, which is the network operator’s perspective. Even though we also consider (partial) track closures, which are typically considered as disruptions, in our examples they will be managed by modifying existing train paths. The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f. [37] and references therein). Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
@@
As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for the efficient suppression of the consequences of disturbances. While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [27], the maximal delay that cannot be avoided [10], or fuel consumption measures  [22]. Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost in real time. The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints [36, 29]. It has an extensive literature, already summarized in a number of reviews, including those by [5], by [26], or by [31].
@@
The maybe most frequently applied modeling strategy is based on alternative graphs [10, 29], resulting in a disjunctive program. Such a model can also be converted into a mixed integer linear program referred to as the big-M model. As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and combinations of these; see the work of [26] for a review of the model and the methods.
@@
An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though the so-arising $0-1$ programs can be large, this method has been applied for both timetabling and dispatching/rescheduling [6, 28, 30]. Meanwhile, $0-1$ programs are suitable as input for certain quantum hardware types. Indeed, job-shop problems addressed on hardware quantum annelaers [39], and even in certain hybrid approaches [25] use similar models, and our previous work [15] which was the first to apply a quantum solver to railway dispatching problems also followed this route. In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP). We introduce a variant of this type of model which maintains a reasonable model size while addressing practically relevant problem instances. In this way, a hybrid classical-quantum integer solver can be applied to solve them, and a comparison with a classical solver is also possible.
@@
In particular, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple-track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We do not include train cancellations in the model. We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model by [17]. Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
@@
We recognize several important scientific gaps. First, existing optimization models do not scale well: they can become huge and inefficient for larger real-life instances. Exploiting the characteristics of a given railway scenario, one can possibly construct simplified models whose size remains small enough to be addressed with a suitable exact or heuristic solver, yet they capture the relevant optimization objectives and constraints. For the particular railway scenario and solver studied in the present paper, no such model is known. Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, no hybrid QA-based models have been used for real-life railway optimization so far, nor in other schedule-based modes, like public transport or air traffic planning/rescheduling. In this paper, we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of an urban railway network.
@@ 2 Quantum annealing
Quantum computing devices became available for practical computational purposes in the last few years. Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics [23, 3]. The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
The class of QUBO problems is well studied and has a number of applications; we refer the reader to a recent monograph edited by [32]. Linear binary problems can also transformed to QUBOs; linear and quadratic constraints can be taken into account with penalties [21]. Using appropriate binary encoding of the variables, integer programming problems can also be treated. Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
Expressing a QUBO as the energy of an Ising system one gets a configuration of pairwise couplings of spins and local fields applied to individual spins, which can be realized in tunable quantum hardware. According to the adiabatic theorem of quantum mechanics [2], tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out. Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems. Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  [11].
In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs. They output a number of configurations typically close to, and possibly including the optimal one.
The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled. The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system. This embedding is a hard problem itself [41]. It often requires to couple multiple physical qubits to represent a logical bit of the problem. As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem can only be set with limited accuracy.
@@
Currently, the applicability of QA technologies is being explored. In the railway context, the first applications of QA can be found in train dispatching/rescheduling [17] and rolling stock planning [4]. In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [15]. A follow-up paper by [17] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research, which is gaining increasing attention, e.g. the recent work of [40].
@@ 3 Problem description
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in additional delays of trains, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. We will minimize the function of these secondary delays.
@@
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as the station, and consider it as a black box, without a detailed layout. We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot. Rolling stock circulation conditions are prescribed to ensure the precedence between the service train and the actual train.
Our model shares features of both microscopic and macroscopic models. At the decision stations, it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting. Meanwhile, from the point of view of the rest of the network, it is a macroscopic model.
@@ 4.1.1 Sets
$\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}$ train pairs possibly competing for an interlocking area of $s$ upon departure
$\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon the departure of $j$ from $s$ while $j^{\prime}$ arrives to $s$ from the direction of $s^{\prime}$.
$\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival from $s^{\prime}$ under special circumstances described in the main text.
$\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival, under special circumstances described in the main text.
@@
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their departure from station $s$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$. For these train pairs, the respective constraints will ensure that they do not pass the same interlocking area at the same time upon departure.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same interlocking area at the same time upon arrival, which is also enforced by the interlocking system. The possibility for M-O between the current and previous decision stations (e.g. using multiple tracks) is to be modeled specifically, hence the distinction between these sets and the $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$-s.
@@
  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction.
  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$.

### v6

@@ Abstract
We address the applicability of a hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver (CQM) as well as with CPLEX, for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments. All the requirements posed by the operator of the network included in the model. The computational results demonstrate the readiness for application and the benefits of quantum-classical hybrid solvers in a realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
@@ 1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policies, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and, especially, railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems. Rail transport, due to its technical and organizational characteristics, is very sensitive to disturbances and disruptions in traffic. These extraordinary events have an impact on railway operations, typically resulting in delays [18]. Examples of such disturbances include late train departures/arrivals, extended dwell times, etc. Their duration can be from several minutes up to hours. A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too [5]. In the present work, we focus on train rescheduling from the network operator’s perspective. Even though we also consider (partial) track closures, which are typically viewed as disruptions, in our examples they will be managed by modifying existing train paths. The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f. [37] and references therein). Ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers and train operators.
@@
A railway network is a complex non-local structure. Hence, modeling a bigger portion of it is necessary for the efficient suppression of the consequences caused by disturbances. Diverse objective functions have been employed for that purpose, such as the (weighted) sum of delays [27], the maximal delay that cannot be avoided [10], or fuel consumption measures  [22]. These large-scale rescheduling problems have to be solved almost in real time, as the time available for making decisions is limited. The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints [36, 29]. This scheduling problem has an extensive literature, already summarized in a number of reviews, including those by [5], by [26], or by [31].
@@
One of the most frequently applied modeling strategies is based on alternative graphs [10, 29], resulting in a disjunctive program. Such a model can also be converted into a mixed integer linear program referred to as the big-M model. As this model often becomes too big to be solved on time, a number of strategies are applied in order to quickly find a suitable solution, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and combinations of these; see the work of [26] for a review of the model and the methods.
@@
An alternative approach to this scheduling problem is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though the so-arising $0-1$ programs can be large, this method has been applied for both timetabling and dispatching/rescheduling [6, 28, 30]. Meanwhile, $0-1$ programs are suitable as input for certain quantum hardware types. Indeed, job-shop problems addressed on hardware quantum annelaers [39], and even in certain hybrid approaches [25] use similar models. Our previous approach [15], which was the first to apply a quantum solver to railway dispatching problems, also followed this route. In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP). We introduce a variant of this type of model which maintains a reasonable model size while addressing practically relevant problem instances. In this way, a hybrid classical-quantum integer solver can be applied to solve them, and a comparison with a classical solver is also possible.
@@
In particular, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple-track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We do not include train cancellations in the model. We apply a new hybrid quantum-classical rescheduling algorithm combining classical heuristics with Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model by [17]. Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
@@
We recognize several important scientific gaps which the present work aims to contribute to. First, existing optimization models do not scale well: they can become huge and inefficient for larger real-life instances. Exploiting the characteristics of a given railway scenario, one can possibly construct simplified models whose size remains small enough to be addressed with a suitable exact or heuristic solver, yet they capture the relevant optimization objectives and constraints. For the particular railway scenario and solver studied in the present paper, no such model is known. Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, to our knowledge, no hybrid QA-based models have been used for real-life railway optimization so far, nor in other schedule-based modes, like public transport or air traffic planning/rescheduling. In this paper, we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of an urban railway network.
@@ 2 Quantum annealing
Quantum computing devices became available for practical computational purposes in the last few years. Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system, well-known in physics [23, 3]. The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables. This is an NP-hard problem in general.
The class of QUBO problems is well studied and has a number of applications; we refer the reader to a recent monograph edited by [32]. Linear and quadratic constrained binary problems can also be transformed to QUBOs: linear and quadratic constraints can be taken into account with penalties [21]. Using appropriate binary encoding of the variables, integer programming problems can be treated in this way as well. Hence, an efficient QUBO solver may be useful in many problems, including the integer linear programs arising in the present contribution.
By expressing a QUBO as the energy of an Ising system, one gets a configuration of pairwise couplings of spins and local fields applied to individual spins, which, in turn, can be realized in tunable quantum hardware. According to the adiabatic theorem of quantum mechanics [2], tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, the optimal configuration can be read out under certain conditions. Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, nevertheless it can be efficient in a variety of problems. Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  [11].
In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided. Hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs. They output a number of configurations typically close to, and possibly including the optimal one.
The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits, where not all pairs of these are coupled. The fixed topology of qubit couplings means that the problem’s graph (defined by the nonzero couplings) has to be embedded as an induced subgraph of the graph describing the topology of the system. This embedding is a computationally hard problem itself [41]. It often requires to couple multiple physical qubits to represent a logical bit of the problem. As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem can only be set with limited accuracy.
@@
Currently, the applicability of QA technologies is being explored in various fields. In the railway context, the first applications of QA can be found in train dispatching/rescheduling [17] and rolling stock planning [4]. In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [15]. A follow-up paper by [17] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research, which is gaining increasing attention, see e.g. the recent work of [40].
@@ 3 Problem description
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in additional delays of trains, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. We will minimize a function of these secondary delays.
@@
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as a station, and consider it as a black box, without a detailed layout. We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot. Rolling stock circulation conditions are prescribed to ensure the precedence between the service train and the actual train.
Our model is mainly macroscopic while it contains some microscopic features, including station technology, track occupancy, rolling stock circulation, and shunting. Concerning switching zones, as no adjusting routes are modeled, independent parts of switching zones can be considered, thus parallel entries and departures can be performed.
@@ 4.1.1 Sets
$\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon departure
$\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon the departure of $j$ from $s$ while $j^{\prime}$ arrives to $s$ from the direction of $s^{\prime}$.
$\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon arrival from $s^{\prime}$ under special circumstances described in the main text.
$\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon arrival, under special circumstances described in the main text.
@@
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their departure from station $s$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$ (given trains’ paths fixed, $s^{\prime}$ is uniquely defined by $j^{\prime}$ and $s$). For these train pairs, the respective constraints will ensure that they do not pass the same switch in an interlocking area at the same time upon departure.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same switch in an interlocking area at the same time upon arrival, which is also enforced by the interlocking system. The possibility for M-O between the current and previous decision stations (e.g. using multiple tracks) is to be modeled specifically, hence the distinction between these sets and the $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$-s.
@@
  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction, and they cannot meet and overtake between the subsequent stations.
  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs ($s$, $s^{\prime}$) that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. (Note that $s^{\prime}$ is uniquely defined by $j$, $j^{\prime}$ and $s$ because the routes of trains are fixed.)